# Automatic Link Selection in Multi-Channel Multiple Access with Link Failures

### v1

@@ III-A Solving Problem (8)
where (a) follows due to Lemma 11, and (b) follows due to Lemma 7. Now, due to Lemma 1, we have $\phi(\boldsymbol{\gamma}^{*})=\phi^{\text{opt}}$. Using Jensen’s inequality, we are done. ∎
IV Discussion
IV-A Adaptiveness
A careful inspection of the proof of Theorem 1 shows that the result holds even if the success probabilities $q_{i,j}$ changed before time $T_{0}$, as long as they remained constant during $[T_{0}:T_{0}+T-1]$. Hence, the adaptiveness is satisfied.
IV-B Enforcing user fairness
Consider the case when $n\geq m$ (the number of users is at least the number of channels). Assume we fix a $\theta\in(0,1/s]$ such that each user is required to transmit at least $\theta$ fraction of the time on average on each channel and no user can transmit more than $(1-(n-m)\theta)$ fraction of the time. This enables faster adaptation to the new optimality point.
To see this, notice that for (P1), we require the additional constraints of
$\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\mathbb{E}\{Y_{i,j}(t)\}\geq\theta.$
for all $i\in[n]$ and $j\in[m]$ and
$\lim\inf_{T\to\infty}\frac{1}{T}\sum_{t=1}^{T}\sum_{j=1}^{m}\mathbb{E}\{Y_{i,j}(t)\}\leq 1-(n-m)\theta,$
for all $i\in[n]$. This will simply transform the constraint $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}$ of (P2) to $\boldsymbol{P}\in\mathcal{S}^{\text{doub}}_{\theta}$. Setting $\varepsilon=\theta$ in Algorithm 1, this will allow us to use $\boldsymbol{P}^{*}$ directly in (1) in Theorem 1, instead of $(1-\varepsilon s)\boldsymbol{P}^{*}+\varepsilon\boldsymbol{1}$, which gives
$\phi^{*}$ $-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}$
$\leq\frac{n}{V}+\frac{9\eta nms^{2}}{2\theta V}(BV+1)^{2}+\frac{\theta nms(BV+1)}{V}+\frac{s}{\eta VT}\log\left(\frac{1}{\theta}\right)+\frac{n(BV+1)^{2}}{2VT}$
$\ \ \ \ +\frac{nB(BV+1)}{T}.$
Now, since we assume $\theta$ to be a constant, using $\eta=\Theta(1/T)$, and $V=\Theta(\sqrt{T})$, we have
$\phi^{*}$ $-\mathbb{E}\left\{\phi\left(\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right)\right\}=O\left(\frac{1}{\sqrt{T}}\right)$
IV-C Case $m=1$
This case can be approached using the work on bandits with vector rewards and concave utility functions [26]. However, the algorithms proposed are not adaptive. It turns out that the algorithm for the general multi-channel case can be simplified for the single-channel case preserving the adaptiveness. In addition, the simplified algorithm has faster convergence. The idea of the algorithm is to find a distribution $\boldsymbol{p}(t)$ over the users in the $(t-1)$-th iteration, after which we sample the user to be assigned on the $t$-th iteration from $\boldsymbol{p}(t)$. Similar to the multi-channel case, we use auxiliary variables $\boldsymbol{\gamma}(t)\in[0,1]^{n}$ and a virtual queue $\boldsymbol{Q}(t)\in\mathbb{R}^{n}_{+}$ for all $t\in\mathbb{N}$. For notational convenience, for each $t\in\mathbb{N}$ and $i\in[n]$, we will use $q_{i},S_{i}(t)$ and $Y_{i}(t)$ instead of $q_{i,1},S_{i,1}(t)$ and $Y_{i,1}(t)$, respectively. The algorithm uses three parameters $\varepsilon\in(0,1/n)$, $V>0$, and $\eta>0$. See Algorithm 2 for details of the implementation.
1 Initialize $\boldsymbol{p}(1)\in\Delta_{n,\varepsilon}$, $\boldsymbol{\gamma}(1)\in[0,1]^{n}$, and the virtual queues $\boldsymbol{Q}\in[0,BV+1]^{n}$.
2 for each iteration $t\in[1:T]$ do
3       Sample $a_{t}\sim\boldsymbol{p}(t)$ and set $\boldsymbol{Y}(t)=\boldsymbol{e}_{a_{t}}$ where $\boldsymbol{e}_{i}$ is the unit vector with $i$-th entry being 1.
4       Receive feedback $\boldsymbol{S}(t)\odot\boldsymbol{Y}(t)$.
5       Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using, $\hat{S}_{i}(t)=\frac{S_{i}(t)Y_{i}(t)}{p_{i}(t)}$ for all $i\in[1:n]$.
6       Find, $\boldsymbol{p}(t+1),\boldsymbol{\gamma}(t+1)$ by solving the problem,
$(\boldsymbol{p}(t+1),\boldsymbol{\gamma}(t+1))=\arg\min_{\boldsymbol{p}\in\Delta_{n,\varepsilon},\boldsymbol{\gamma}\in[0,1]^{n}}\Big{[}-V\phi(\boldsymbol{\gamma})$
$+\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}-\hat{S}_{i}(t)p_{i}]+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\Big{]},$ (29)
where $D_{\text{KL}}$ is the KL-divergence.
7       Update the virtual queues,
$\boldsymbol{Q}(t+1)=[\boldsymbol{Q}(t)+\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)]_{+},$ (30)
where $\boldsymbol{X}(t)=\boldsymbol{Y}(t)\odot\boldsymbol{S}(t)$.
8
Algorithm 2 Single Channel Adpative MAC
Most of the results required for the error analysis are borrowed from the multi-channel case. We only prove the results that are unique to this case. We first focus on solving the intermediate problem (6). Notice that the problem can be separated into the problem of finding optimal $\boldsymbol{\gamma}(t+1)$ and $\boldsymbol{p}(t+1)$.
@@ 7       Update the virtual queues,
Finding $\boldsymbol{\gamma}(t+1)$: This is the same problem as in the multi-channel case.
Finding $\boldsymbol{p}(t+1)$: This is (P3) with $\boldsymbol{x}=\eta\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)$, and $\boldsymbol{y}=\boldsymbol{p}(t)$.
Similar to Lemma 5, we have Pinsker’s inequality for KL divergence.
Lemma 12 (Pinsker’s inequality)
For $\boldsymbol{x},\boldsymbol{y}\in\Delta_{n}$, we have that,
$D_{\text{KL}}(\boldsymbol{x}\lVert\boldsymbol{y})\geq\frac{1}{2}\lVert\boldsymbol{x}-\boldsymbol{y}\rVert_{1}^{2}\geq\frac{1}{2}\lVert\boldsymbol{x}-\boldsymbol{y}\rVert^{2},$
We also have the following lemma similar to Lemma 6.
Lemma 13
We have $D_{\text{KL}}(\boldsymbol{x}\lVert\boldsymbol{y})\leq\log\left(\frac{1}{\varepsilon}\right)$, for all $\boldsymbol{x}\in\Delta_{n}$ where $\boldsymbol{y}\in\Delta_{n,\varepsilon}$.
Notice that since the queue updates and the problem to solve $\boldsymbol{\gamma}(t+1)$ are the same as in the multi-channel case, we have the same queue bound.
Lemma 14
We have that for all $t\in\{1,2,,\dots\}$ and $i\in\{1,2\dots,n\}$, $Q_{i}(t)\leq BV+1$.
Define the drift $\Delta(t)$ as,
@@
We have that,
-- $\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q_{i}(t)p_{i}(t)\right\}$
$\ \ \ \ \leq\frac{\eta n}{2\varepsilon}(BV+1)^{2}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}-\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\},$
where $\hat{S}_{i}(t)$ is defined in line 5 of Algorithm 2.
@@
Notice that,
$\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\}$
$=\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)[p_{i}(t+1)-p_{i}(t)]\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$
$\leq_{(a)}\frac{\eta}{2}\mathbb{E}\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}+\frac{1}{2\eta}\mathbb{E}\left\{\lVert\boldsymbol{p}(t+1)-\boldsymbol{p}(t)\rVert^{2}\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$
$\leq_{(b)}\frac{\eta}{2}\mathbb{E}\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}+\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}$ (32)
where (a) follows from $\frac{1}{2\eta}\lVert\boldsymbol{a}\rVert^{2}+\frac{\eta}{2}\lVert\boldsymbol{b}\rVert^{2}\geq\sum_{i=1}^{n}a_{i}b_{i}$ for $n$-dimensional vectors $\boldsymbol{a,b}$, (b) follows from Lemma 12. Define the history $\mathcal{H}(t)$ similar to multi-channel case (see, (23)). Notice that $\boldsymbol{p}(\tau),\boldsymbol{Q}(\tau),\boldsymbol{\gamma}(\tau)$ for $\tau\in[t]$ are $\mathcal{H}(t)$-measurable. Now, we handle the terms of (16) separately. Notice that,
$\mathbb{E}$$\left\{\lVert\boldsymbol{Q}(t)\odot\boldsymbol{\hat{S}}(t)\rVert^{2}\right\}=\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\hat{S}_{a_{t}}^{2}(t)\right\}$
$=\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\frac{S_{a_{t}}(t)}{p_{a_{t}}(t)}\hat{S}_{a_{t}}(t)\right\}\leq_{(a)}\frac{1}{\varepsilon}\mathbb{E}\left\{Q^{2}_{a_{t}}(t)\hat{S}_{a_{t}}(t)\right\}=\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i}(t)}{p_{i}(t)}Y_{i}(t)\right\}$
$=\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\mathbb{E}\left\{Q^{2}_{i}(t)\frac{S_{i}(t)}{p_{i}(t)}Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}=_{(b)}\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\frac{Q_{i}(t)^{2}}{p_{i}(t)}\mathbb{E}\left\{S_{i}(t)Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}$
$=_{(c)}\frac{1}{\varepsilon}\sum_{i=1}^{n}\mathbb{E}\left\{\frac{Q_{i}(t)^{2}}{p_{i}(t)}q_{i}\mathbb{E}\left\{Y_{i}(t)\Bigg{|}\mathcal{H}(t)\right\}\right\}=\frac{1}{\varepsilon}\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q^{2}_{i}(t)\right\}\leq_{(d)}\frac{1}{\varepsilon}\mathbb{E}\{\lVert\boldsymbol{Q}(t)\rVert^{2}\}$
$\leq\frac{n}{\varepsilon}(BV+1)^{2}$
where $a_{t}$ is defined in line 3 of Algorithm 2, (a) follows since $S_{a_{t}}(t)\leq 1$ and $p_{a_{t}}(t)\geq\varepsilon$, (b) follows since $\boldsymbol{Q}(t)$ and $\boldsymbol{p}(t)$ are $\mathcal{H}(t)$-measurable, (c) follows since $\boldsymbol{S}(t)$ is independent of $\boldsymbol{Y}(t)$ and $\mathcal{H}(t)$, (d) follows since $q_{i}\leq 1$, and the last inequality follows due to Lemma 14. Also,
$\sum_{i=1}^{n}$ $\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t)\right\}=\mathbb{E}\left\{Q_{a_{t}}(t)\frac{S_{a_{t}}(t)}{p_{a_{t}}(t)}p_{a_{t}}(t)\right\}=\sum_{i=1}^{n}q_{i}\mathbb{E}\left\{Q_{i}(t)p_{i}(t)\right\}$
Combining everything and substituting in (16), we are done. ∎
@@
We have that for any $T_{0}\in\mathbb{N}$, $\boldsymbol{\gamma}\in[0,1]^{n}$, and $\boldsymbol{p}\in\Delta_{n,\varepsilon}$,
@@
$\leq nT+\frac{\eta nT}{2\varepsilon}(BV+1)^{2}+\sum_{t=T_{0}}^{T+T_{0}-1}\sum_{i=1}^{n}(\gamma_{i}-q_{i}p_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$
@@
Adding $-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 15, we have that,
@@
$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}\mathbb{E}\{Q_{i}(t)p_{i}(t)\}$
$\leq_{(a)}n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{\eta n}{2\varepsilon}(BV+1)^{2}$
$\ \ \ \ \ +\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}-\sum_{i=1}^{n}\mathbb{E}\left\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\right\},$ (33)
where (a) follows by Lemma 16. Now notice that combining the decision (6) with Lemma 4, we have that for any $\boldsymbol{\gamma}\in[0,1]^{n}$ and $\boldsymbol{p}\in\Delta_{n,\varepsilon}$,
$-V\phi(\boldsymbol{\gamma}(t+1))+\sum_{i=1}^{n}\gamma_{i}(t+1)Q_{i}(t)+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))-\sum_{i=1}^{n}Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)$
$\ \ \ \ \leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}Q_{i}(t)[\gamma_{i}-\hat{S}_{i}(t)p_{i}]+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))-\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1)).$
Taking expectations of the above, we have that,
$-V$ $\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ \ -\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}$
$\leq-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}\}+\frac{1}{\eta}\mathbb{E}\left\{D(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}$
$=-V\phi(\boldsymbol{\gamma})+\sum_{i=1}^{n}\gamma_{i}\mathbb{E}\{Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}p_{i}\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\right\}$
$\ \ \ \ -\frac{1}{\eta}\mathbb{E}\left\{D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t+1))\right\}.$
Substituting the above in (17), we have that,

### v2

@@ V-B Analysis of the Algorithm
where (a) follows due to Lemma 11, and (b) follows due to Corollary 1.1. Now, due to Lemma 1, we have $\phi(\boldsymbol{\gamma}^{*})=\phi^{*}\geq\phi^{\text{opt}}$. Combining with Jensen’s inequality, we are done.
VI Special Cases
In this section, we introduce several interesting cases of the general setting. The first is an adaptive algorithm for the single-channel case ($m=1$). This algorithm has both closed form solutions for iterations and fast convergence. Then we introduce a UCB-based non-adaptive algorithm with efficient implementation. Both algorithms will be used for comparisons. Finally, we consider the single channel ($m=1$) special case with $\phi(\boldsymbol{x})=\min\{x_{1},x_{2},\dots,x_{n}\}$. In this setting, a simple mechanism achieves convergence to the optimal value while adapting to channel variations, without requiring explicit estimation.
VI-A Single Channel Algorithm
The single channel case simplifies due to the absence of matching constraints. In particular, in each iteration, we assign a user to the channel, and the assigned user in the $t$-th time slot is sampled from $\boldsymbol{p}(t)\in\Delta^{n}$. For this section, we use the notation $S_{i}(t)=S_{i,1}(t)$, $Y_{i}(t)=Y_{i,1}(t)$. Algorithm 3 summarizes the steps.
1Initialize $\boldsymbol{p}(1)=\frac{1}{n}\boldsymbol{1}$ and the virtual queues $\boldsymbol{Q}(1)=\boldsymbol{0}\in\mathbb{R}^{n}$.
2 for each time slot $t\in\mathbb{N}$ do
3   Sample $l_{t}\in[n]$ from $\boldsymbol{p}(t)$, set $\boldsymbol{Y}(t)$ as the $l_{t}$-th unit vector, and receive feedback.
4   Compute the estimator $\boldsymbol{\hat{S}}(t)$ for $\boldsymbol{S}(t)$ using $\hat{S}_{i}(t)=\frac{S_{i}(t)Y_{i}(t)}{p_{i}(t)}$ for all $i\in[n]$
5   Find
$\boldsymbol{p}(t+1)=\arg\min_{\boldsymbol{p}\in\Delta^{n}_{\varepsilon}}\Big[-\sum_{i=1}^{n}Q_{i}(t)\hat{S}_{i}(t)p_{i}+\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}\lVert\boldsymbol{p}(t))\Big],$ (65)
6   Update $\boldsymbol{Q}(t+1)=\left[\boldsymbol{Q}(t)+\boldsymbol{\gamma}(t+1)-\boldsymbol{X}(t)\right]_{+}$.
Algorithm 3 Single-Channel Adaptive MAC (Parameters $V>0$, $\varepsilon\in(0,1/n])$, and $\eta>0$)
We first state the performance bound of the algorithm.
Theorem 3
Consider $T,T_{0}\in\mathbb{N}$ and suppose that the success probabilities $q_{i,j}$ remain fixed on the interval $T_{0},\,T_{0}+T-1]$. For any parameters $\varepsilon\in(0,1/n)$, $\eta,V>0$, we have that,
$\leq\frac{n}{V}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta V}+\frac{\varepsilon n^{2}(BV+1)}{V}+\frac{1}{\eta VT}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2VT}+\frac{nB(BV+1)}{T}$
In particular, using $\eta=\Theta(\sqrt{\log(T)}/T)$, $\varepsilon=\Theta(\sqrt{\log(T)}/\sqrt{T})<1$, and $V=\Theta(T^{1/2})$, we have
$\phi^{\text{opt}}-\phi\left(\mathbb{E}\left\{\frac{1}{T}\sum_{t=T_{0}}^{T+T_{0}-1}\boldsymbol{X}(t)\right\}\right)=O\left(\sqrt{\frac{\log(T)}{T}}\right).$
VI-A1 Solving Problem (5)
@@
Notice that finding $\boldsymbol{\gamma}(t+1)$ is the same problem as in the multi-channel case. In addition, the problem to solve for $\boldsymbol{p}(t+1)$ reduces to the problem (P3) discussed in Appendix D.
VI-A2 Deterministic Problem
Consider the problem
$\text{s.t. }\boldsymbol{p}\in\Delta_{n}$ (67)
$\ \ \ \ \ \boldsymbol{\gamma}\in[0,1]^{n}$ (68)
$\ \ \ \ \ q_{i}p_{i}\geq\gamma_{i}\ \forall i\in\{1,\dots,n\},$ (69)
Using an argument similar to Lemma 1, it can be shown that the optimal objective value of the above problem is at least $\phi^{\text{opt}}$. For this section, let $\phi^{*}$ denote the optimal objective value of the above problem.
Let $(\boldsymbol{p}^{*},\boldsymbol{\gamma}^{*})$ denote an optimal solution of the above problem. If $q_{1},q_{2},\dots,q_{n}$ were known, the policy of sampling a user independently in each time slot using the distribution $\boldsymbol{p}^{*}$ achieves an objective value of $\phi^{*}$ for (P1). The subsequent analysis shows that even when $q_{1},q_{2},\dots,q_{n}$ are unknown, we can achieve an objective value arbitrarily close to $\phi^{*}$ for (P1), thereby showing $\phi^{*}=\phi^{\text{opt}}$.
VI-A3 Analysis of the Algorithm
Most of the results required for the error analysis are borrowed from the multi-channel case. In particular, notice that the deterministic queue bound Corollary 1.1 holds in this case.
Define the drift $\Delta(t)$ as
@@ VI-A3 Analysis of the Algorithm
For any $t\in[T_{0}:T+T_{0}-1]$, we have
- $\sum_{i=1}^{n}q_{i}\mathbb{E}\{p_{i}(t)Q_{i}(t)\}$
$\leq-\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}+\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}$ (71)
where $\hat{S}_{i}(t)$ is defined in line 4 of Algorithm 3.
@@
Using Lemma 5, we have
$\sum_{i=1}^{n}Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)$
$\leq_{(a)}\frac{1}{\eta}\Bigg[D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\ln\left(\sum_{i=1}^{n}{p}_{i}(t)e^{\eta Q_{i}(t)\hat{S}_{i}(t)}\right)\Bigg]$
$\leq_{(b)}\frac{1}{\eta}\Bigg[D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\ln\left(\sum_{i=1}^{n}{p}_{i}(t)\left[1+\eta Q_{i}(t)\hat{S}_{i}(t)+\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{(BV+1)^{2}}Q^{2}_{i}(t)\hat{S}^{2}_{i}(t)\right]\right)\Bigg]$
$=\frac{1}{\eta}\Bigg[D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\ln\left(1+\sum_{i=1}^{n}\eta{p}_{i}(t)Q_{i}(t)\hat{S}_{i}(t)+\sum_{i=1}^{n}\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{(BV+1)^{2}}p_{i}(t)Q^{2}_{i}(t)\hat{S}^{2}_{i}(t)\right)\Bigg]$
$\leq_{(c)}\frac{1}{\eta}\Bigg[D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\sum_{i=1}^{n}\eta{p}_{i}(t)Q_{i}(t)\hat{S}_{i}(t)+\sum_{i=1}^{n}\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{(BV+1)^{2}}p_{i}(t)Q^{2}_{i}(t)\hat{S}^{2}_{i}(t)\Bigg]$
$=_{(d)}\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\sum_{i=1}^{n}{p}_{i}(t)Q_{i}(t)\hat{S}_{i}(t)+\sum_{i=1}^{n}\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}S_{i}(t)Y_{i}(t)\hat{S}_{i}(t)$
$=_{(e)}\frac{1}{\eta}D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))+\sum_{i=1}^{n}{p}_{i}(t)Q_{i}(t)\hat{S}_{i}(t)+\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}\sum_{i=1}^{n}\hat{S}_{i}(t)$ (72)
where (a) follows from Lemma 5, (b) follows combining Corollary 1.1 and $\hat{S}_{i}(t)\leq 1/\varepsilon$ with $e^{x}\leq 1+x+(e^{c}-c-1)x^{2}/c^{2}$ for $x\in[0,c]$ for any $c>0$, (c) follows from $\ln(1+x)\leq x$, (d) follows from the definition of $\hat{S}_{i}(t)$, and the last inequality follows since $S_{i}(t),Y_{i}(t)\in\{0,1\}$. Taking the expected value conditioned on $\mathcal{H}(t)$, followed by taking unconditional expectations, we have
$\sum_{i=1}^{n}\mathbb{E}\{Q_{i}(t)\hat{S}_{i}(t)p_{i}(t+1)\}$
$\leq\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\}+\sum_{i=1}^{n}q_{i}\mathbb{E}\{p_{i}(t)Q_{i}(t)\}+\frac{\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}\sum_{i=1}^{n}q_{i}$
$\leq\frac{1}{\eta}\mathbb{E}\{D_{\text{KL}}(\boldsymbol{p}(t+1)\lVert\boldsymbol{p}(t))\}+\sum_{i=1}^{n}q_{i}\mathbb{E}\{p_{i}(t)Q_{i}(t)\}+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}}{\eta}$ (73)
where the last inequality follows since $q_{i}\in[0,1]$. Rearranging, we are done. ∎
@@
We have that for any $T_{0}\in\mathbb{N}$, $\boldsymbol{\gamma}\in[0,1]^{n}$, and $\boldsymbol{p}\in\Delta^{n}_{\varepsilon}$,
@@
$\leq nT+\frac{n\varepsilon^{2}e^{\frac{\eta(BV+1)}{\varepsilon}}T}{\eta}+\sum_{t=T_{0}}^{T_{0}+T-1}\sum_{i=1}^{n}(\gamma_{i}-p_{i}q_{i})\mathbb{E}\{Q_{i}(t)\}+\frac{1}{\eta}\log\left(\frac{1}{\varepsilon}\right)+\frac{n(BV+1)^{2}}{2}$ (74)
@@
Adding $V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}$ to the result of Lemma 16, we have that,
@@
$\leq n-V\mathbb{E}\{\phi(\boldsymbol{\gamma}(t+1))\}+\sum_{i=1}^{n}\mathbb{E}\{\gamma_{i}(t+1)Q_{i}(t)\}-\sum_{i=1}^{n}q_{i}\mathbb{E}\{Q_{i}(t)p_{i}(t)\}$ (75)
Using Lemma 17 on the above, we have